OpenStudy (anonymous):

Please show me how to solve 5-6x => 2x + 21

OpenStudy (anonymous):

5-6x-2x-2=0 3-8x=0 3=8x x=3/8

OpenStudy (anonymous):

I got x <= -2

OpenStudy (anonymous):

few screwing you were right

OpenStudy (anonymous):

Well there's a first for everything ;)

OpenStudy (anonymous):

i am typo,...

OpenStudy (anonymous):

I'm sure everyone's familiar with the process (David sh just made a simple error), but here's the working:\[\Large \begin{array}{l} {\rm{5}} - {\rm{6x }} \ge {\rm{ 2x }} + {\rm{ 21}}\\ - 6x - 2x \ge 21 - 5\\ - 8x \ge 16\\ - x \ge \frac{{16}}{8}\\ - x \ge 2\\ x \le - 2 \end{array}\]

OpenStudy (anonymous):

yup thanks. because my teacher did this problem, and she put x < -2, so she must've made a mistake.

OpenStudy (anonymous):

Ah it's easy to do, especially if you've been writing out numbers all day.

OpenStudy (anonymous):

lol i figured